OpenStudy (anonymous):

What is the value of x in the rhombus below? If necessary, round your answer to two decimal places.

jhonyy9 (jhonyy9):

???

jhonyy9 (jhonyy9):

AC = 3x AB = x+2 yes ?

OpenStudy (anonymous):

Yea

OpenStudy (anonymous):

I have an attachment

jhonyy9 (jhonyy9):

yes i got it so hope do you know that the diagonales of a rhombus bisect yes ?

OpenStudy (anonymous):

Yea

jhonyy9 (jhonyy9):

let the bisect intersection point O so than AO = 3x/2 yes ?

OpenStudy (anonymous):

Yea

jhonyy9 (jhonyy9):

do you have any choises ?

OpenStudy (anonymous):

No, it's just a solver

jhonyy9 (jhonyy9):

can you calculi OB ?

OpenStudy (anonymous):

Its my last question

OpenStudy (anonymous):

OB theres no OB

OpenStudy (anonymous):

You mean DB?

jhonyy9 (jhonyy9):

no OB what mean DB halfed because O is the intersection point ok ?

OpenStudy (anonymous):

No.. I got it.. its 3x+x+2=90

OpenStudy (anonymous):

so 4x=88 x=22

OpenStudy (anonymous):

Right

jhonyy9 (jhonyy9):

i dont know ?

jhonyy9 (jhonyy9):

why 3x+x+2 = 90 ?

OpenStudy (anonymous):

Because this is geometry

OpenStudy (anonymous):

they are compelementary

OpenStudy (anonymous):

Thanks for help, I got it! :D